OpenStudy (anonymous):

Find the area of the region between the two curves y=3x-x^2 and y = 2x^3-x^2-5x

OpenStudy (anonymous):

3x-x^2=2x^3-x^2-5x to find limits

OpenStudy (eyust707):

I usually graph these first just to see what were dealing with: http://www.wolframalpha.com/input/?i=graph+y%3D3x-x^2+and+y+%3D+2x^3-x^2-5x

OpenStudy (eyust707):

then id use an iterated integral: \[\int\limits_{0}^{2} \int\limits_{2x^3-x^2-5x}^{3x-x^2} dydx\]

OpenStudy (eyust707):

i think... lol

OpenStudy (anonymous):

wow wtf is that :D

OpenStudy (eyust707):

is this for calc 2 or calc 3?

OpenStudy (anonymous):

i would integrate (3x-x^2-(2x^3-x^2-5x))dx

OpenStudy (anonymous):

How to Integrate something like that?

OpenStudy (anonymous):

it's the same

OpenStudy (eyust707):

you learn how in calc3

OpenStudy (anonymous):

im not from us

OpenStudy (eyust707):

is just the inside then the outside

OpenStudy (anonymous):

i don't know your calcs

OpenStudy (anonymous):

if you solve the inner integral it's the same what tomas got

OpenStudy (anonymous):

wow I love the multiple integral

OpenStudy (anonymous):

mine looks so much easier :P

OpenStudy (anonymous):

of course it is it's just that eyust looks more advanced lol

OpenStudy (eyust707):

lol i always make s**t more complicated dooohhhhh

OpenStudy (anonymous):

i got 8 as the answer